OpenStudy (anonymous):

what is 4 over 16 in simplest form? a- 1 over 8 b- 1 over 4 c- 1 over 2 d- 2 over 8

OpenStudy (anonymous):

@bibby

OpenStudy (esshotwired):

divide both both sides by 4, which is the GCF. it is B

OpenStudy (anonymous):

ok thanks wait ok

OpenStudy (esshotwired):

wait for what?

OpenStudy (anonymous):

Drew notices there are twenty-one girls in a class of thirty students. What is the simplest form of twenty-one thirtieths? three-tenths three-fifths three-sevenths seven-tenths

OpenStudy (anonymous):

@esshotwired

OpenStudy (esshotwired):

\[\frac{ 21\div3 }{ 30\div3 }\] 3 is the GCF so divide both sides by 3. the answer is D 7/10

OpenStudy (anonymous):

Which fraction is equivalent to eight over twenty? four over five two over ten two over five one over four

OpenStudy (anonymous):

@esshotwired

OpenStudy (esshotwired):

\[\frac{ 8 }{ 20 }\] 1. find the simplest form \[\frac{ 2 }{ 5 }\] is the simplest form and it is an answer choice so the answer is C